# Security modelling for cyber-physical systems: A systematic literature review

### v1

@@ 4.3 Categorisation and classification
We categorised the selected articles into three clusters according to the following criteria:
  • Cluster 1 (C1): Article focuses on threat modelling method(s)
  • Cluster 2 (C2): Article focuses on attack modelling method(s)
  • Cluster 3 (C3): Article is a systematic literature review (SLR)
Additionally, we classified the articles according to the following criteria:
@@ 4.5 Data collection
We studied the articles closely to understand the background, approaches, and results from the various research. Importantly, we sought to identify gaps in the research in terms of security modelling in the context of CPS cyber attacks. We documented our notes from the review in a shared repository (Huang (2024)) for future reference.
@@ 5 Results
The search was conducted on the following digital libraries in January 2024: IEEE Computer Society Digital Library, Science Direct, ACM Digital Library and Scopus. Following the search, we had 438 articles in total. We performed a rigorous quality assessment, where we first excluded duplicate articles, then we identified and excluded irrelevant articles based on their abstracts. We then performed full-reading of about 27 articles that remained and excluded articles that were not relevant to any of our research questions. Finally, we had 15 articles subject to further analysis (Figure 5).
Figure 5: Search results (January 2024)
@@
Table 2: Categorisation of articles
Reference
Title
Threat Modelling (C1)
Attack Modelling (C2)
SLRs (C3)
Ayrour et al. (2018)
Modelling cyber attacks: a survey study
$\bullet$
Fernandez (2016)
Threat modeling in cyber-physical systems
$\bullet$
Khalil et al. (2022)
Threat Modeling of Cyber-Physical Systems-A Case Study of a Microgrid System
$\bullet$
Khalil et al. (2023)
Threat modeling of industrial control systems: A systematic literature review
$\bullet$
Khan et al. (2017)
STRIDE-based threat modeling for cyber-physical systems
$\bullet$
Kim et al. (2022)
STRIDE‐based threat modeling and DREAD evaluation for the distributed control system in the oil refinery
$\bullet$
Kumar, R., et al. (2022)
APT attacks on industrial control systems: A tale of three incidents
$\bullet$
Martins et al. (2015)
Towards a systematic threat modeling approach for cyber-physical systems
$\bullet$
Mekdad et al. (2021)
A threat model method for ICS malware: the TRISIS case
$\bullet$
Neubert and Vielhauer (2020)
Kill chain attack modelling for hidden channel attack scenarios in industrial control systems
$\bullet$
Paudel et al. (2017)
Attack models for advanced persistent threats in smart grid wide area monitoring
$\bullet$
Tatam et al. (2021)
A review of threat modelling approaches for APT-style attacks
$\bullet$
Xiong and Lagerström (2019)
Threat modeling – A systematic literature review
$\bullet$
Yang and Zhang (2023)
From Tactics to Techniques: A Systematic Attack Modeling for Advanced Persistent Threats in Industrial Control Systems
$\bullet$
Zalewski et al. (2013)
Threat modeling for security assessment in cyberphysical systems
$\bullet$
@@
2014201520162017201820192020202120222023111122223333
Figure 6: Number of articles per year
Next, we used citation metrics on Google Scholar to identify articles that had more than 100 citations. As of January 2024, only two of the 15 selected articles had more than 100 citations (Table 3). The top cited article was on STRIDE-based threat modelling for cyber-physical systems by Khan et al. (2017), followed by the SLR on threat modelling by Xiong and Lagerström (2019). This observation, together with that from the number of articles published yearly, suggested that this field of research is still relatively new.
Table 3: Top-cited articles
@@
Three of the articles (Ayrour et al. (2018); Kumar et al. (2022); Mekdad et al. (2021)) addressed all the research questions. The remaining articles addressed some, but not all the research questions. 12 of the 15 articles were relevant to RQ1, and only two of the seven C1 articles were relevant to RQ1.1 (28.57%), one was relevant to RQ1.2 (14.29%), and three (42.86%) were relevant to RQ2. On the other hand, three of the four C2 articles were relevant to RQ1.1 (75%), four were relevant to RQ1.2 (100%) and two were relevant to RQ2 (50%). Lastly, three of the four C3 articles were relevant to RQ1.1 (75%), one was relevant to RQ1.2 (25%), and four were relevant to RQ2.
@@ 6.1.1 RQ1
To answer the first research question, “what cyber-physical system (CPS) threat or attack models have been adopted in existing literature?”, we aimed to identify the threat and attack models that were adopted in the selected articles. This was not straightforward as authors appeared to have varying definitions of threat and attack modelling methods, and SLRs in particular, categorised both as threat modelling methods.
For example, Khalil et al. (2023) proposed “attack-centric threat modelling” in their ICS threat modelling taxonomy, and distinguished the term from attack modelling, stating that the latter focused solely on the attacker’s behaviour and not the defender’s perspective. In addition, Tatam et al. (2021) categorised models that are typically used to analyse cyber attacks, such as the diamond intrusion analysis model (Caltagirone et al., 2013), attack trees (Saini et al. (2008); Schneier (1999)), the kill chain (Hutchins et al. (2011)), and the MITRE ATT&CK framework (Strom et al. (2018)) as threat models.
As such, we used the clusters as a basis for identifying threat and attack modelling methods described in the articles (Table 5), since the categorisation was based on research objectives and approach. We found that this approach produced a reasonable result as the modelling methods appeared to be categorised correctly. One exception was the article by Mekdad et al. (2021) where the diamond intrusion analysis model may be more accurately categorised as an attack modelling method.
@@
Reference Cluster
Threat Modelling Method(s)
Attack Modelling Method(s)
Fernandez (2016) C1
Misuse patterns
Khalil et al. (2022) C1
STRIDE
Khan et al. (2017) C1
STRIDE
Kim et al. (2022) C1
STRIDE, DREAD
Martins et al. (2015) C1
Generic Modeling Environment (GME)
Mekdad et al. (2021) C1
Diamond Model of Intrusion Analysis
Zalewski et al. (2013) C1
STRIDE
Kumar et al. (2022) C2
Attack tree
Neubert and Vielhauer (2020) C2
Kill chain
Paudel et al. (2017) C2
Attack tree
Yang and Zhang (2023) C2
Abstract APT attack model
@@
On the second research question, “how do the proposed threat and attack models align with cybersecurity in various phases of the cyber-physical system lifecycle?”, we found that threat modelling was conducted in the early stages of the system development cycle (Khalil et al. (2022)), or more specifically, during system design and validation phases (Khan et al. (2017); Kim et al. (2022); Martins et al. (2015)). Interestingly, unlike threat modelling, none of the C2 articles suggested or mentioned which phase of the system lifecycle attack modelling should be performed in. A possible reason is that attack modelling is seen as a complement to threat modelling methods (Khalil et al. (2022); Kumar et al. (2022)).
@@
On the third research question, “in what manner do the proposed threat and attack models adjust to evolving attacker tactics, techniques, and procedures over time?”, we found that there were different views on incorporating attacker tactics, techniques, and procedures (TTPs) into threat modelling frameworks. For example, Khalil et al. (2022) viewed that attack categories in their proposed attack taxonomy should be abstract and should not include details of the attack techniques such as those in MITRE ATT&CK for Industrial Control Systems (Alexander et al. (2020)). In contrast, Fernandez (2016) proposed a threat model using misuse patterns which may include attacker TTPs to represent CPS threats.
Likewise, several articles made references to attacker TTPs, or similar attacker behavioural characteristics in their analysis. Mekdad et al. (2021) combined the Diamond Model of intrusion analysis (Caltagirone et al. (2013)) and the ICS kill chain (Assante and Lee (2015)) to generate the ICS threat model. Attack trees (Schneier (1999)) were used to model CPS attacks (Kumar et al. (2022); Paudel et al. (2017)), while Neubert and Vielhauer (2020) presented how the Lockheed Martin Cyber Kill Chain (Hutchins et al. (2011)) can be used to model hidden channel attack scenarios in ICS. Yang and Zhang (2023) proposed an abstract APT attack model that can recognise and correlate attacker TTPs to identify and mitigate complex attacks in ICS, highlighting that the need to consider the range of TTPs used by sophisticated adversaries to represent real threats and reflect the evolving threat landscape.
Importantly, we noted that while several articles considered the need for threat or attack models to account for changes over time, they did not include attacker TTPs as one of those changes. For example, Almohri et al. (2017) recognised that the threat model of a medical CPS is related to its corresponding trust model, and the trust model may include time-bound elements such as temporarily-trusted individuals or accounts. Additionally, Khalil et al. (2022) noted that systems may change over time and information initially excluded from threat analysis should be included in future threat models to account for system modifications.
@@ 6.1.4 RQ2
For our final research question, “what research gaps have been identified in the methods used for modelling threats and attacks in cyber-physical systems (CPS)?”, we made several observations from the literature.
@@
Security models mostly focus on IT systems. Firstly, threat and attack models in the literature mostly focus on IT systems, even in the context of CPS security. For example, the STRIDE model developed by Microsoft was used in several articles to identify CPS threats, notwithstanding that the model is more commonly used in the context of software development and system security. Using STRIDE in CPS threat modelling (Abuabed et al. (2023); Khan et al. (2017); Kim et al. (2022)) has several limitations which were not obvious in the literature.
@@
Specifically, the STRIDE approach relies on Data Flow Diagrams (DFDs) to identify system entities and trust boundaries, which has several shortcomings related to inadequate representation of security concepts, data elements, abstraction levels, and deployment information (Sion et al. (2020)). Moreover, it is sometimes not practical or even possible to construct an accurate DFD in CPS. There are various reasons for this: a CPS is a system-of-systems that includes both cyber and physical components; availability, accuracy and integrity of data flows between cyber and physical components cannot be assumed; and most importantly, the consequences of CPS compromise extend beyond those specified in the STRIDE model, e.g. physical changes or effects.
@@
This limitation may be addressed by leveraging diverse, multi-disciplinary expertise from physical security and domain specialists to augment the DFD by capturing interdependencies and interactions between cyber and physical components, and not be limited to IT systems; and by continuously refining the DFD based on real-time data and feedback from system operations and cybersecurity events detected in the CPS environment over time.
@@
Relatedly, the literature refers to CPS use-cases in the context of industrial control systems (ICS), smart grids, medical CPS (MCPS), supervisory control and data acquisition (SCADA) systems, or modern vehicles. We expect that CPS use-cases will increasingly include new scenarios involving technologies such as robotics, drones, EV charging infrastructure and smart buildings. Conventional threat modelling approaches, primarily tailored for IT systems, may encounter heightened challenges when addressing cybersecurity threats in these emerging domains.
@@
Attack scenarios are not realistic. Secondly, most threat and attack modelling methods in the literature made assumptions based on simplistic cybersecurity attack scenarios, and did not consider the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS. The exceptions were: Fernandez (2016) who proposed a patterns-based threat model that considers multiple vulnerabilities exploited in different parts of a distributed system, and Zografopoulos et al. (2021) who analysed Cyber-Physical Energy Systems (CPES) security from a multi-layered attack perspective.
@@
Differentiation between IT and CPS cybersecurity breaches. Thirdly, most of the literature did not differentiate cybersecurity breaches in CPS from those in IT systems in that they are not a binary (all or nothing) event. Instead of abrupt, catastrophic system failures, CPS cybersecurity breaches often result in gradual degradation of system services before the CPS fails eventually (Zalewski et al. (2013)). This is analogous to a system fault in an autonomous vehicle, the vehicle does not come to a stop immediately—which would injure the passengers—but instead gradually slows to a safe stop. The implication to CPS security is that from a cyber defence perspective, there is an opportunity to remediate intrusions through cyber or physical interventions (Ertaul and Mousa (2018)) or to perform dynamic reconfiguration and self-healing tasks (Zografopoulos et al. (2021)) before the CPS fails completely.
@@
Consequence-driven considerations. CPS increasingly face cybersecurity threats that can endanger the safety, reliability and resilience of critical infrastructure. With the convergence of cybersecurity, safety, and reliability in CPS, it would be beneficial to adopt a consequence-driven and cyber-informed approach (Freeman et al. (2016)) towards threat and attack modelling, and align with system development processes in engineering domains such as the V-model that is commonly used in the automotive industry (Bolz et al. (2020)), as well as the failure mode and effect analysis (FMEA) and fault-tree analysis (FTA) techniques used in safety and reliability engineering domains (Kriaa et al. (2015)). Doing so will avoid making threat or attack modelling a one-time event that assumes static cybersecurity threats and vulnerabilities, resulting in inherent security weaknesses that are “baked” into proprietary CPS systems and software designs.
@@
Distinguishing between CPS threat and attack modelling. Lastly, we found that while the literature used “threat modelling” and “attack modelling” interchangeably, both terms usually referred to threat modelling and not the latter. On the other hand, some authors promoted definitions of threat modelling from literature and distinguished threat from attack modelling (Khalil et al. (2023); Xiong and Lagerström (2019)), while others proposed security modelling frameworks that are based on threat models (Burmester et al. (2012)). Based on the literature, we analysed the various characteristics and inter-relationships between threat and attack modelling methods and these are aligned with our proposed unified security modelling framework for CPS as described in Section 2.4.

### v2

@@ 4.3 Categorisation and classification
We categorised the selected articles into three distinct clusters, based on their primary focus and content:
  • Cluster 1 (C1): Articles focused on threat modelling method(s)
    Articles are placed in this cluster if their primary focus is on threat modelling, defined as the process of identifying potential threats and vulnerabilities early in the system lifecycle to inform the design of mitigation strategies. These articles predominantly discuss frameworks, tools, techniques, or methodologies specifically designed for threat modelling.
  • Cluster 2 (C2): Articles focused on attack modelling method(s)
    Articles are included in this cluster if they primarily focus on attack modelling, which involves analysing attacker tactics, techniques, and procedures (TTPs) to inform defence strategies during the system’s operational phase. These articles are centered on understanding the behavior, capabilities, and goals of attackers and often provide detailed descriptions of attacker models, simulation methods, or real-world case studies of cyber attacks.
  • Cluster 3 (C3): Articles that are systematic literature reviews (SLRs)
    Articles are categorised in this cluster if they are systematic literature reviews that provide a comprehensive review of existing threat and/or attack modelling methods. These articles systematically analyse and synthesise findings from a range of studies, offering insights into the state of research, common practices, gaps, and future directions.
Articles were categorised based on a thorough analysis of their content, including their primary objectives, methodologies, and the specific aspects of threat or attack modelling they emphasise. Where articles covered both threat and attack modelling, they were classified according to the dominant theme or focus, based on the extent and depth of discussion related to either method.
Additionally, articles were categorised according to the following criteria:
@@ 4.5 Data collection
We studied the articles closely to understand the background, approaches, and results from the various research. Importantly, we sought to identify gaps in the research in terms of security modelling in the context of CPS cyber attacks. We documented our notes from the review in a public repository [41] for future reference.
@@ 5 Results
The search was conducted on the following digital libraries in August 2024: IEEE Computer Society Digital Library, Science Direct, ACM Digital Library and Scopus. Following the search, we had 366 articles in total. We performed a rigorous quality assessment, where we first excluded duplicate articles, then we identified and excluded irrelevant articles based on their abstracts. We then performed full-reading of about 32 articles that remained and excluded articles that were not relevant to any of our research questions. Finally, we had 15 articles subject to further analysis (Figure 5).
Figure 5: Literature survey search results (August 2024)
@@
Table 1: Categorisation of CPS cybersecurity articles
Reference Title Threat Modelling (C1) Attack Modelling (C2) SLRs (C3)
Ayrour et al. (2018) Modelling cyber attacks: a survey study $\bullet$
Fernandez (2016) Threat modeling in cyber-physical systems $\bullet$
Khalil et al. (2022) Threat Modeling of Cyber-Physical Systems-A Case Study of a Microgrid System $\bullet$
Khalil et al. (2023) Threat modeling of industrial control systems: A systematic literature review $\bullet$
Khan et al. (2017) STRIDE-based threat modeling for cyber-physical systems $\bullet$
Kim et al. (2022) STRIDE‐based threat modeling and DREAD evaluation for the distributed control system in the oil refinery $\bullet$
Kumar, R., et al. (2022) APT attacks on industrial control systems: A tale of three incidents $\bullet$
Martins et al. (2015) Towards a systematic threat modeling approach for cyber-physical systems $\bullet$
Mekdad et al. (2021) A threat model method for ICS malware: the TRISIS case $\bullet$
Neubert and Vielhauer (2020) Kill chain attack modelling for hidden channel attack scenarios in industrial control systems $\bullet$
Paudel et al. (2017) Attack models for advanced persistent threats in smart grid wide area monitoring $\bullet$
Tatam et al. (2021) A review of threat modelling approaches for APT-style attacks $\bullet$
Xiong and Lagerström (2019) Threat modeling – A systematic literature review $\bullet$
Yang and Zhang (2023) From Tactics to Techniques: A Systematic Attack Modeling for Advanced Persistent Threats in Industrial Control Systems $\bullet$
Zalewski et al. (2013) Threat modeling for security assessment in cyberphysical systems $\bullet$
@@
Figure 6: Number of CPS cybersecurity articles per year
Next, we used citation metrics on Google Scholar to identify articles that had more than 100 citations. As of January 2024, only two of the 15 selected articles had more than 100 citations (Table 2). The top cited article was on STRIDE-based threat modelling for cyber-physical systems by [42], followed by the SLR on threat modelling by [8]. This observation, together with that from the number of articles published yearly, suggested that this field of research is still relatively new.
Table 2: Top-cited CPS cybersecurity articles
@@
Three of the articles [43] [3] [44] addressed all the research questions. The remaining articles addressed some, but not all the research questions. 12 of the 15 articles were relevant to RQ1, and only two of the seven C1 articles were relevant to RQ1.1 (28.57%), one was relevant to RQ1.2 (14.29%), and three (42.86%) were relevant to RQ2. On the other hand, three of the four C2 articles were relevant to RQ1.1 (75%), four were relevant to RQ1.2 (100%) and two were relevant to RQ2 (50%). Lastly, three of the four C3 articles were relevant to RQ1.1 (75%), one was relevant to RQ1.2 (25%), and four were relevant to RQ2.
@@ 6.1.1 RQ1
To address the first research question, “What CPS threat or attack models have been adopted in existing literature?”, we examined the threat and attack modelling methods described in the articles, excluding the SLRs. SLRs were excluded because they primarily focus on summarising and synthesising existing research rather than presenting original threat or attack modelling methods. Therefore, they provide an overview of methodologies rather than specific models. The methods we identified include: the Diamond Intrusion Analysis model [23], attack trees [10] [11], the Lockheed Martin cyber kill chain [21], and the MITRE ATT&CK framework [19], as detailed in Table 4.
@@
Reference Cluster Threat and Attack Modelling Method(s)
Fernandez (2016) C1 Misuse patterns
Khalil et al. (2022) C1 STRIDE
Khan et al. (2017) C1 STRIDE
Kim et al. (2022) C1 STRIDE, DREAD
Martins et al. (2015) C1 Generic Modeling Environment (GME)
Mekdad et al. (2021) C1 Diamond Model of Intrusion Analysis
Zalewski et al. (2013) C1 STRIDE
Kumar et al. (2022) C2 Attack tree
Neubert and Vielhauer (2020) C2 Kill chain
Paudel et al. (2017) C2 Attack tree
Yang and Zhang (2023) C2 Abstract APT attack model
@@
On the second research question, “how do the proposed threat and attack models align with cybersecurity in various phases of the CPS lifecycle?”, we found that threat modelling was conducted in the early stages of the system development cycle [45], or more specifically, during system design and validation phases [42] [46] [47]. Interestingly, unlike threat modelling, none of the C2 articles suggested or mentioned which phase of the system lifecycle attack modelling should be performed in. A possible reason is that attack modelling is seen as a complement to threat modelling methods [45] [3].
@@
On the third research question, “in what manner do the proposed threat and attack models adjust to evolving attacker tactics, techniques, and procedures over time?”, we found differing views on how to incorporate attacker tactics, techniques, and procedures (TTPs) into threat modelling frameworks. For example, Khalid et al. [45] argued that attack categories in their proposed taxonomy should be abstract and should not include specific details of attack techniques, such as those outlined in MITRE ATT&CK for Industrial Control Systems [20]. In contrast, Fernandez et al. [48] proposed a threat model using misuse patterns that may include attacker TTPs to represent CPS threats. These differing approaches highlight the challenge of balancing model specificity with adaptability to evolving threats.
Several studies have attempted to incorporate attacker TTPs or similar behavioural characteristics into their models. For instance, Mekdad et al. [44] combined the Diamond Model of Intrusion Analysis [23] and the ICS kill chain [49] to develop a dynamic ICS threat model. Similarly, attack trees [11] have been employed to model CPS attacks [3] [14], while Neuber et al. [50] demonstrated how the Lockheed Martin Cyber Kill Chain [21] could be used to model hidden channel attack scenarios in ICS. Yang et al. [13] proposed an abstract APT attack model capable of recognising and correlating attacker TTPs to identify and mitigate complex attacks in ICS. This model emphasises the need to consider a broad range of TTPs to reflect the evolving threat landscape accurately.
However, our analysis revealed that while some models incorporate elements of adaptability to evolving TTPs, there is a significant gap in addressing how these models dynamically adjust to the continuously changing tactics of sophisticated adversaries. For example, while models like those proposed by Mekdad et al. and Yang et al. suggest mechanisms for integrating new TTPs, they do not specify processes for continuously updating or validating these models against emerging threats.
Moreover, while several articles acknowledged the need for models to account for changes over time, they did not explicitly address attacker TTPs as a component of these changes. For instance, Almohri et al. [36] noted that the threat model of a medical CPS might change with its corresponding trust model, incorporating time-bound elements such as temporarily trusted individuals or accounts. Similarly, Khalil et al. [45] highlighted that systems may evolve over time, necessitating the inclusion of previously excluded information in future threat models to account for system modifications.
@@
This gap suggests a need for more dynamic and resilient security models that can adapt to both the evolving threat landscape and the changing nature of CPS environments.
@@ 6.1.4 RQ2
For our final research question, “what research gaps have been identified in the methods used for modelling threats and attacks in CPS?”, we made several observations from the literature.
@@
Security models mostly focus on IT systems. Firstly, threat and attack models in the literature mostly focus on IT systems, even in the context of CPS security. For example, the STRIDE model developed by Microsoft was used in several articles to identify CPS threats, notwithstanding that the model is more commonly used in the context of software development and system security. Using STRIDE in CPS threat modelling [34] [42] [46] has several limitations which were not obvious in the literature.
@@
Specifically, the STRIDE approach relies on Data Flow Diagrams (DFDs) to identify system entities and trust boundaries, which has several shortcomings related to inadequate representation of security concepts, data elements, abstraction levels, and deployment information [51]. Moreover, it is sometimes not practical or even possible to construct an accurate DFD in CPS. There are various reasons for this. A CPS is a system-of-systems that includes both cyber and physical components. Availability, accuracy and integrity of data flows between cyber and physical components cannot be assumed. Most importantly, the consequences of CPS compromise extend beyond those specified in the STRIDE model, e.g. physical changes or effects.
@@
This limitation may be addressed by leveraging diverse, multi-disciplinary expertise from physical security and domain specialists to augment the DFD by capturing interdependencies and interactions between cyber and physical components, and not be limited to IT systems. This limitation may also be addressed by continuously refining the DFD based on real-time data and feedback from system operations and cybersecurity events detected in the CPS environment over time.
@@
Relatedly, the literature refers to CPS use-cases in the context of ICS, smart grids, medical CPS (MCPS), SCADA systems, or modern vehicles. We expect that CPS use-cases will increasingly include new scenarios involving technologies such as robotics, drones, EV charging infrastructure and smart buildings. Conventional threat modelling approaches, primarily tailored for IT systems, may encounter heightened challenges when addressing cybersecurity threats in these emerging domains.
@@
Attack scenarios are not realistic. Secondly, most threat and attack modelling methods in the literature made assumptions based on simplistic cybersecurity attack scenarios, and did not consider the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS. The exceptions were: Fernandez et al. [48] who proposed a patterns-based threat model that considers multiple vulnerabilities exploited in different parts of a distributed system, and Zografopoulos et al. [33] who analysed Cyber-Physical Energy Systems (CPES) security from a multi-layered attack perspective.
@@
Differentiation between IT and CPS cybersecurity breaches. Thirdly, most of the literature did not differentiate cybersecurity breaches in CPS from those in IT systems in that they are not a binary (all or nothing) event. Instead of abrupt, catastrophic system failures, CPS cybersecurity breaches often result in gradual degradation of system services before the CPS fails eventually [52]. This is analogous to a system fault in an autonomous vehicle, the vehicle does not come to a stop immediately—which would injure the passengers—but instead gradually slows to a safe stop. The implication to CPS security is that from a cyber defence perspective, there is an opportunity to remediate intrusions through cyber or physical interventions [24] or to perform dynamic reconfiguration and self-healing tasks [33] before the CPS fails completely.
@@
Consequence-driven considerations. CPS increasingly face cybersecurity threats that can endanger the safety, reliability and resilience of critical infrastructure. With the convergence of cybersecurity, safety, and reliability in CPS, it would be beneficial to adopt a consequence-driven and cyber-informed approach [53] towards threat and attack modelling, and align with system development processes in engineering domains such as the V-model that is commonly used in the automotive industry [54], as well as the failure mode and effect analysis (FMEA) and fault-tree analysis (FTA) techniques used in safety and reliability engineering domains [16]. Doing so will avoid making threat or attack modelling a one-time event that assumes static cybersecurity threats and vulnerabilities, resulting in inherent security weaknesses that are “baked” into proprietary CPS systems and software designs.
@@
Distinguishing between CPS threat and attack modelling. Lastly, we found that while the literature used “threat modelling” and “attack modelling” interchangeably, both terms usually referred to threat modelling and not the latter. On the other hand, some authors promoted definitions of threat modelling from literature and distinguished threat from attack modelling [28] [8], while others proposed security modelling frameworks that are based on threat models [26]. In this paper, we provide explicit definitions for both threat and attack modelling to clarify their distinct roles within cybersecurity (see Section 1). Based on these definitions and the literature, we analysed the various characteristics and inter-relationships between threat and attack modelling methods and these are aligned with our proposed unified security modelling framework for CPS as described in Section 2.4.